OpenStudy (anonymous):

Part 1: Explain, in complete sentences, which method you would use to solve the following system of equations. (1 point) Part 2: Explain why you chose that method. (1 point) Part 3: Provide the solution to the system. (2 points) x – 3y + 4z = –9 2x + y + 3z = 6 x – 4y + 3z = –9

OpenStudy (anonymous):

use lpp

jhonyy9 (jhonyy9):

x-3y+4z=-9 2x+y+3z=6 x-4y+3z=-9 solve by elimination method x-3y+4z=-9 multiplie by -2 and will get -2x +6y -8z=18 2x +y +3z = 6 ------------------ add term by term 0 7y -5z = 24 2x+y+3z=6 x -4y+3z=-9 multiplie by -2 and will get -2x+8y-6z=18 2x +y +3z=6 -2x+8y-6z=18 --------------- add term by term 0 9y -3z =24 than you have got first 7y-5z=24 multiplie first equation by 3 to eliminate adding the term of z 9y-3z=24 multiplie second equation by -5 21y-15z=72 -45y+15z=-120 ----------------- adding will get -24y 0 = -48 -24y=-48 y= -48/(-24) = 2 y=2 7y-5z=24 7*2-5z=24 14-5z=24 -5z=10 z= 10/(-5) = -2 z = -2 y = 2 2x+y+3z=6 2x +2 +3(-2)=6 2x +2 -6 =6 2x +4=6 2x= 2 x=2/2 x =1 y = 2 z = -2 hope so much that this will help you good luck bye